using law of sines, find all missing measurements in the triangle KZM:
angle K= 75 degrees
angle Z= 53 degrees
and side KZ=26

Answer by stanbon(75887)

find all missing measurements in the triangle KZM:
---
using law of sines, find all missing measurements in the triangle KZM:
angle K= 75 degrees
angle Z= 53 degrees
and side KZ=26
---------------------------
The side opposite angle K is ZM
The side opposite angle Z is KM
The angle opposite KZ is M
----------------------------------
Angle M = 180-(75+53)
Angle M = 52 degrees
------
Law of Sines:
MZ/sin(75) = KZ/sin(52)
---
MZ = (sin(75))[26/sin(52)]
---
MZ = 31.87
=====================
MK/sin(53)=26/sin(52)
---
MK = (sin(53))[26/sin(52)]
---
MK = 26.35
====================
Cheers,
Stan H.